# OneDiff: A Generalist Model for Image Difference Captioning

### v1

@@ Abstract
In computer vision, Image Difference Captioning (IDC) is crucial for accurately describing variations between closely related images. Traditional IDC methods often rely on specialist models, which restrict their applicability across varied contexts. This paper introduces the OneDiff model, a novel generalist approach that utilizes a robust vision-language model architecture, integrating a siamese image encoder with a Visual Delta Module. This innovative configuration allows for the precise detection and articulation of fine-grained differences between image pairs. OneDiff is trained through a dual-phase strategy, encompassing Coupled Sample Training and multi-task learning across a diverse array of data types, supported by our newly developed DiffCap Dataset. This dataset merges real-world and synthetic data, enhancing the training process and bolstering the model’s robustness. Extensive testing on diverse IDC benchmarks, such as Spot-the-Diff, CLEVR-Change, and Birds-to-Words, shows that OneDiff consistently outperforms existing state-of-the-art models in accuracy and adaptability, achieving improvements of up to 85% CIDEr points in average. By setting a new benchmark in IDC, OneDiff paves the way for more versatile and effective applications in detecting and describing visual differences. The code, models, and data will be made publicly available.
[…]
1476
@@ 1 Introduction
Humans possess a remarkable ability to detect differences within visual scenes, a skill critical for myriad real-world applications such as change detection in surveillance footage, anomaly detection in medical images, and more. This human capability has inspired the development of Image Difference Captioning (IDC), a task in computer vision aimed at algorithmically describing discrepancies between two similar images. IDC is pivotal for precise difference detection required in fields like medical diagnostics, environmental monitoring, and manufacturing quality control. Additionally, IDC holds potential for aiding in tasks like distinguishing closely related species, describing lesions in medical images, and monitoring changes in media assets.
@@
To effectively train our generalist IDC model, we create DiffCap Dataset, an open-domain, large-scale dataset comprising three distinct types of IDC data: aggregated task-specific collections and two forms of synthetic data that we developed. First, we aggregate existing IDC datasets to form a solid base of varied real-world scenarios. Second, we expand our dataset with open-domain real image pairs, identified through image library searches and equipped with GPT-assisted autogenerated captions to increase coverage and context variability. Finally, we enhance DiffCap with synthetic images from InstructPix2Pix dataset, which are specifically designed to include subtle and complex changes. These synthetic images are paired with change summaries generated by GPT-3, addressing the dual challenges of training data scarcity and limited manipulation variety in conventional IDC datasets. This strategic combination allows our model to adapt and perform robustly across a wide array of scenarios.
@@ 3.1 Task-Specific Data Collection
We firstly curate data from existing datasets that focus on identifying differences between image pairs, as is shown in Figure 1. These include benchmarks like Spot-the-Diff [9] and CLEVR-Change, which are directly compatible with our IDC tasks. Additionally, the Birds-to-Words dataset, which is tailored for describing differences between birds, and the Image-Editing-Request dataset, which provides instructions for editing images, are also included. The NLVR2 dataset, known for its dual-image inputs, further enhances the model’s ability to capture intricate details in image pairs.
@@ 3.2 Automatic Data Construction with Real Images
We implement a retrieval-based approach to collect similar images from a large image database, LAION-5B, with source images from MSCOCO [14] and Flickr30K [18] datasets. The image similarity is computed as the cosine similarity between two CLIP image features. We filter out image pairs with a similarity score below 0.75 due to high disparity and above 0.94 due to near identity. Additionally, we exclude image pairs with highly similar textual information to ensure the visibility of granular visual differences during annotation.
@@
Following the collection of similar image pairs, we gather necessary textual information for annotation by downstream large language models. Original images sourced from COCO [14] and Flickr30K [18] include manually annotated captions, while retrieved images from LAION-5B feature associated HTML alt-text. Despite the filtering for visual-textual similarity, the alt-text quality from LAION-5B often remains low for use as captions. We employed BLIP-2 to generate new, higher-quality captions for the retrieved images. These captions, along with the original HTML alt-text, are fed into ChatGPT to generate descriptive captions that articulate the differences between the image pairs.
@@ 3.3 Automatic Data Construction with Synthetic Images
While image pairs retrieved by our methods generally display high-level semantic differences, they still exhibit significant low-level visual discrepancies. To better capture these fine-grained details, we have included synthetic images with subtle differences in our dataset.
@@
Our method incorporates the generated paired image editing examples from the InstructPix2Pix framework. InstructPix2Pix uses StableDiffusion in combination with Prompt-to-Prompt to generate pairs of synthetic images from pairs of closely related captions, ensuring each synthetic image pair contains subtle, yet significant, differences relevant to IDC tasks.
@@
We then annotate the difference captions for these synthetic image pairs. Given that the synthetic image pairs are derived from text prompts that guide the Stable Diffusion and Prompt-to-Prompt processes, these prompts inherently contain the modifications made to the original image. However, traditional image editing annotations often focus only on the outcome, missing the context of the original content. To bridge this gap, we employ ChatGPT to generate difference captions that comprehensively describe the changes between the original and edited images. We prompt ChatGPT with the original image prompt, the edited image prompt, and the specific editing instruction, which produces a difference caption. For instance, if the original prompt describes "a serene lake scene" and the editing prompt modifies it to "a serene lake scene with a kayak," the direct instruction might simply be "add a kayak." However, for our IDC tasks, it is more informative to generate a difference caption like "a kayak has been added to the serene lake scene," which explicitly points out the addition within the unchanged context of the lake scene.
@@ 3.4 Dataset Statistics
We develop automatic pipelines to generate image difference captioning data from both realistic and synthetic images for our DiffCap dataset. We have selected 53k pairs of real images, each annotated with difference captions averaging 12.2 words in length. Our synthetic part includes 139k pairs of images and corresponding annotations with an average length of 15.0 words. As supplement, we collect and sample data from existing benchmarks with dual-image inputs, forming a data mixture of 170k image pairs. Aggregating data from above three parts, we have compiled our dataset DiffCap which totals 363k entries detailed in Table 1.
Table 1: Statistics of our DiffCap dataset. ‘Real" refers to our generated data from real images, and ‘Synthetic" refers to our generated data from synthetic images.
Component Sub-Component #Image-Pairs #Annotations
Collected Spot-the-Diff [9] 11K 21K
CLEVR-Change 8K 44K
Image-Editing-Request 3.4K 4.2K
Birds-to-Words 14K 14K
NLVR2 86K 86K
Generated Real 54K 54K
Synthetic 139K 139K
Total 316K 363K
@@ 4.1 Model Architecture
We employ a Transformer for interaction between vision encoders of two images. Specifically, a fixed number of learnable embeddings as so called delta tokens, which are utilized and input to the Transformer consisting of self-attention layers and cross-attention layers alternately, with image hidden states extracted from the encoder layers at equal intervals. These embeddings function as queries for image features in cross-attention layers, thus updated separately and simultaneously with pairs of image information.
@@ 5.1 Implementation Details
The architecture incorporates a vision encoder, CLIP-ViT-L/14, a two-layer multilayer perceptron (MLP) projector, and the large language model (LLM) Vicuna-7B. We utilize the penultimate layer outputs of the CLIP vision encoder for image feature extraction. This encoder consists of 24 layers. From each 336×336 image, we extracted 576 patches of size 24×24. Each patch possesses a dimensionality of 2048, which are subsequently mapped to 4096 to align within the LLM. In the Visual Delta Module (VDM) module, the input learnable delta tokens are initialized using a Gaussian distribution with a mean of 0 and a standard deviation of 0.02. This module is inherited from pretrained BERT-base-uncased [5]. It consists of six layers featuring alternating self-attention and cross-attention blocks. We select image features from the 8th, 16th, and 24th layers of the vision encoder, facilitating interaction among them through delta tokens within the cross-attention blocks.
@@
Our model is trained using the PyTorch framework on 8 NVIDIA A800 (80G) GPUs. During the pre-training phase, the initial learning rate is set at 1e-3 for MLP and 3e-4 for VDM, with a 3% warm-up using a cosine decay schedule. Each GPU processes batches of 32. In the subsequent Supervised Fine-Tuning (SFT) phase, both the learning rate and batch size are reduced to 2e-5 and 8 per GPU, respectively. During pre-training, only the VDM and MLP modules are unfrozen, whereas in the SFT phase, a full fine-tuning of all parameters is implemented.
@@
In the pre-training phase, we employ LLaVA 558k [15] and 279k pairwise image captions generated by CST strategy to construct the pre-training data. In the SFT phase, we initially collect the currently available data, encompassing the training splits of three aforementioned benchmarks, CLEVR-Change [10] and NLVR2 [22], termed the Coll set. This set, along with both real and synthetic data obtained through our data generation pipeline, forms 335k image-text pairs designed for difference caption training, i.e. DiffCap. Furthermore, to preserve the foundational visua-language comprehension capabilities of the model, we also incorporate LLaVA-665k, resulting in a comprehensive dataset of 1M pairs for instruction fine-tuning. Specifically, CLEVR-Change [10], constructed automatically using the CLEVR engine, catalogs geometric object transformations against a plain background. The "after" images capture changes related to color, texture, addition, removal, and movement, alongside minimal viewpoint alterations. NLVR2 [22], a dataset designed for reasoning from natural language descriptions of paired photos, evaluates if visual data supports the descriptions. We randomly select a subset of NLVR2, retaining its original reasoning framework, and enhance it by adding instructions to the textual descriptions to generate data conducive to instruction-following tasks. For descriptions previously marked as "True", we have refined them to include a new instruction: "Make a statement about these two images."
Table 2: Comparison to SOTA on Image-Editing-Request.
@@
OneDiff (ours) 24.6 103.9 24.1 52.2
Table 3: Comparison to SOTA on Spot-the-Diff.
@@
Table 4: Comparison to SOTA on Birds-to-Words.
@@
OneDiff (ours) 25.8 28.0 15.6 49.1
5.2 Evaluation Benchmarks and Metrics
We conduct experiments primarily on three benchmarks: Spot-the-Diff [9], Image-Editing-Request [24], and Birds-to-Words [6], which are described in further detail below. Consistent with prior research, we evaluate our models using BLEU (B) [16], METEOR (M) [4], CIDEr (C) [29], and ROUGE-L (R) [13] as metrics. Specifically, Spot-the-Diff [9] dataset is the inaugural dataset designed for the task of image difference captioning. It comprises 13,192 precisely aligned images sourced from surveillance cameras, characterized by a consistent viewpoint. Each image pair is annotated with a concise description of all significant differences, averaging 1.86 captions per pair. Image-Editing-Request [24] dataset, sourced from 3,939 real image pairs derived from online image-editing forums, includes 5,695 editing directives. Each pair in the training subset is associated with a single directive, whereas those in the validation and test subsets each contain three directives. Typically, the modified objects in these images are subtle and indistinct. Birds-to-Words [6] dataset documents the intricate distinctions among various bird species gathered in natural settings. The dataset contains 3,386 image pairs; each pair is associated with an average of 4.74 detailed captions. Annotators focus solely on the visual differences between birds without referencing specific bird categories or the environmental background. This approach ensures that annotations are both precise and relevant to the observed phenotypic variations.
@@ 5.3 Performance Comparison
As shown in Table 2, we evaluate the performance of our model on Image-Editing-Request. OneDiff demonstrates significant superiority in this benchmark. Specifically, compared to the SOTA models, the improvement of our OneDiff on CIDEr is nearly doubled, i.e. 211% compare with SCORER+CBR. We then compare the performance on Spot-the-Diff. As detailed in Table 3, OneDiff achieves significant improvements across BLEU, CIDEr and METEOR compared with SOTAs. Finally, we analyze the performance disparities on the Birds-to-Words. Consistent with the previous two benchmarks, our model demonstrated a significant lead. As shown in the Table 4.
Table 5: Ablation of model design, including Coupled Sample Training (CST) strategy and Visual Delta Module (VDM). We conduct ablation analysis on three aforementioned benchmarks based on BLEU-4 (B), CIDEr (C), METEOR (M), and ROUGE-L (R).
@@ B C M R B C M R B C M R
× × 26.0 89.3 21.8 49.3 10.0 46.1 13.2 32.8 24.1 18.5 24.8 46.1
× 24.0 96.1 23.3 50.6 10.7 48.6 14.1 33.6 25.5 23.0 25.7 46.4
× 26.1 91.8 23.0 49.8 10.8 46.1 13.6 32.4 24.1 17.7 25.3 45.9
24.6 103.9 24.1 52.2 10.6 47.8 13.9 33.0 26.0 25.0 26.2 47.7
@@
Coll 24.0 94.1 23.3 50.4 8.7 40.6 14.1 31.6 24.5 19.3 25.7 46.4
Coll+Real 24.8 100.1 24.0 51.1 9.7 41.3 14.3 32.8 25.4 20.0 25.8 46.6
Coll+Syn 25.5 95.1 24.0 51.5 9.0 42.0 14.3 32.8 25.5 21.8 25.5 46.7
Coll+Real+Syn 25.8 107.0 24.2 53.6 9.0 42.7 13.2 32.0 25.9 21.1 25.6 46.7

### v3

@@ Abstract
In computer vision, Image Difference Captioning (IDC) is crucial for accurately describing variations between closely related images. Traditional IDC methods often rely on specialist models, which restrict their applicability across varied contexts. This paper introduces the OneDiff model, a novel generalist approach that utilizes a robust vision-language model architecture, integrating a siamese image encoder with a Visual Delta Module. This innovative configuration allows for the precise detection and articulation of fine-grained differences between image pairs. OneDiff is trained through a dual-phase strategy, encompassing Coupled Sample Training and multi-task learning across a diverse array of data types, supported by our newly developed DiffCap Dataset. This dataset merges real-world and synthetic data, enhancing the training process and bolstering the model’s robustness. Extensive testing on diverse IDC benchmarks, such as Spot-the-Diff, Image-Editing-Request, and Birds-to-Words, shows that OneDiff consistently outperforms existing state-of-the-art models in accuracy and adaptability, achieving improvements of up to 97% CIDEr points in average. By setting a new benchmark in IDC, OneDiff paves the way for more versatile and effective applications in detecting and describing visual differences. The code, models, and data will be made publicly available.
Keywords:
Vision-language Model Image Difference Captioning Automatic Data Generation
@@ 1 Introduction
Humans possess a remarkable ability to detect differences within visual scenes, a skill critical for myriad real-world applications such as change detection in surveillance footage, anomaly detection in medical images, and more. This human capability has inspired the development of Image Difference Captioning (IDC), a task in computer vision aimed at algorithmically describing discrepancies between two similar images. IDC is pivotal for precise difference detection required in fields like medical diagnostics, environmental monitoring, and manufacturing quality control. Additionally, IDC holds potential for aiding in tasks like distinguishing closely related species, describing lesions in medical images, monitoring changes in media assets, and differentiating actions in videos [28, 29].
@@
Figure 1: Illustration of Image Difference Captioning tasks.
@@
To effectively train our generalist IDC model, we create DiffCap Dataset, an open-domain, large-scale dataset comprising three distinct types of IDC data: aggregated task-specific collections and two forms of synthetic data that we developed. First, we aggregate existing IDC datasets to form a solid base of varied real-world scenarios. Second, we expand our dataset with open-domain real image pairs, identified through image library searches and equipped with GPT-assisted autogenerated captions to increase coverage and context variability. Finally, we enhance DiffCap with synthetic images from InstructPix2Pix [5] dataset, which are specifically designed to include subtle and complex changes. These synthetic images are paired with change summaries generated by GPT-3 [6], addressing the dual challenges of training data scarcity and limited manipulation variety in conventional IDC datasets. This strategic combination allows our model to adapt and perform robustly across a wide array of scenarios.
@@ 3.1 Task-Specific Data Collection
We firstly curate data from existing datasets that focus on identifying differences between image pairs, as is shown in Fig. 2. These include benchmarks like Spot-the-Diff [12] and CLEVR-Change, which are directly compatible with our IDC tasks. Additionally, the Birds-to-Words dataset, which is tailored for describing differences between birds, and the Image-Editing-Request dataset, which provides instructions for editing images, are also included. The NLVR2 dataset, known for its dual-image inputs, further enhances the model’s ability to capture intricate details in image pairs.
@@ 3.2 Automatic Data Construction with Real Images
We implement a retrieval-based approach to collect similar images from a large image database, LAION-5B, with source images from MSCOCO [17] and Flickr30K [21] datasets. Based on the cosine similarity between two CLIP image features. We filter out image pairs with a similarity score below 0.75 due to high disparity and above 0.94 due to near identity. In addition, we exclude image pairs with highly similar textual information to ensure the visibility of granular visual differences during annotation.
@@
Following the collection of similar image pairs, we gather necessary textual information for annotation by downstream large language models. Original images sourced from COCO [17] and Flickr30K [21] include manually annotated captions, while retrieved images from LAION-5B feature associated HTML alt-text. Despite the filtering for visual-textual similarity, the alt-text quality from LAION-5B often remains low for use as captions. We employed BLIP-2 to generate higher-quality captions for the retrieved images. To alleviate possible hallucinations introduced by BLIP-2, We calculate the image-text similarity between retrieved images and their generated captions using CLIP, and filter for data with scores over 40. These captions, along with the HTML alt-texts, are fed into ChatGPT to generate descriptive captions that articulate the differences between image pairs.
@@ 3.3 Automatic Data Construction with Synthetic Images
While image pairs retrieved by our methods generally display high-level semantic differences, they still exhibit significant low-level visual discrepancies. To better capture these fine-grained details, we have included synthetic images with subtle differences in our dataset to improve generalization capabilities by data synergy.
@@
Our method incorporates the generated paired image editing examples from the InstructPix2Pix [5] framework. InstructPix2Pix uses StableDiffusion in combination with Prompt-to-Prompt to generate pairs of synthetic images from pairs of closely related captions, ensuring each synthetic image pair contains subtle, yet significant, differences relevant to IDC tasks.
@@
Given that the synthetic image pairs are derived from text prompts guiding the Stable Diffusion and Prompt-to-Prompt processes, these prompts inherently contain the modifications made to the original image. However, traditional image editing annotations often focus only on the outcome, missing the context of the original content. To bridge this gap, we employ ChatGPT to generate difference captions that comprehensively describe the changes between the original and edited images. We prompt ChatGPT with the original image prompt, the edited image prompt, and the specific editing instruction to produce a difference caption. For instance, if the original prompt describes "a serene lake scene" and the editing prompt modifies it to "a serene lake scene with a kayak," the direct instruction might simply be "add a kayak."
@@ 3.4 Dataset Statistics
We develop automatic pipelines to generate image difference captioning data from both realistic and synthetic images for our DiffCap dataset. We have selected 53k pairs of real images, each annotated with difference captions averaging 12.2 words in length. Our synthetic part includes 139k pairs of images and corresponding annotations with an average length of 15.0 words. As supplement, we collect and sample data from existing benchmarks with dual-image inputs, forming a data mixture of 170k image pairs. Aggregating data from above three parts, we have compiled our dataset DiffCap which totals 363k entries detailed in our supplementary materials.
@@ 4.1 Model Architecture
We employ a Transformer for interaction between vision encoders of two images. Specifically, a fixed number of learnable embeddings as so called delta tokens, which are utilized and input to the Transformer consisting of self-attention layers and cross-attention layers alternately, with image hidden states extracted from the encoder layers at equal intervals. These embeddings function as queries for image features in cross-attention layers, thus updated separately and simultaneously with pairs of image information. Our method can be conveniently adapted for models projecting encoded image features into the textual space.
@@ 5.1 Implementation Details
The architecture incorporates a vision encoder, CLIP-ViT-L/14, a two-layer multilayer perceptron (MLP) projector, and the large language model (LLM) Vicuna-7B [7]. We utilize the penultimate layer outputs of the CLIP vision encoder for image feature extraction. This encoder consists of 24 layers. From each 336×336 image, we extracted 576 patches of size 24×24. Each patch possesses a dimensionality of 2048, which are subsequently mapped to 4096 to align within the LLM. In the Visual Delta Module (VDM) module, the input learnable delta tokens are initialized using a Gaussian distribution with a mean of 0 and a standard deviation of 0.02. This module is inherited from pretrained BERT-base-uncased [8]. It consists of six layers featuring alternating self-attention and cross-attention blocks. We select image features from the 8th, 16th, and 24th layers of the vision encoder, facilitating interaction among them through delta tokens within the cross-attention blocks.
@@
Our model is trained on 8 NVIDIA A800 (80G) GPUs for 10 hours (Vicuna) or 4 hours (Qwen2 [37]). During the pre-training phase, the initial learning rate is set at 1e-3 for MLP and 3e-4 for VDM, with a 3% warm-up using a cosine decay schedule. Each GPU processes batches of 32. In the subsequent Supervised Fine-Tuning (SFT) phase, both the learning rate and batch size are reduced to 2e-5 and 8 per GPU, respectively. During pre-training, only the VDM and MLP modules are unfrozen, whereas in the SFT phase, a full fine-tuning of all parameters is implemented.
@@
In the pre-training phase, we employ LLaVA 558k [18] and 279k pairwise image captions generated by CST strategy to construct the pre-training data. In the SFT phase, we initially collect the currently available data, encompassing the training splits of three aforementioned benchmarks, CLEVR-Change [13] and NLVR2 [25], termed the Coll set. This set, along with both real and synthetic data obtained through our data generation pipeline, forms 335k image-text pairs designed for difference caption training, i.e. DiffCap. Furthermore, to preserve the foundational visua-language comprehension capabilities of the model, we also incorporate LLaVA-665k, resulting in a comprehensive dataset of 1M pairs for instruction fine-tuning. Specifically, CLEVR-Change [13], constructed automatically using the CLEVR engine, catalogs geometric object transformations against a plain background. The "after" images capture changes related to color, texture, addition, removal, and movement, alongside minimal viewpoint alterations. NLVR2 [25] evaluates whether paired visual data supports the statements. We randomly select a subset of NLVR2, retaining its original reasoning framework, and enhance it by adding instructions to generate data conducive to instruction-following tasks. For "True" statements, we refine them to include a new instruction: "Make a statement about these two images."
5.2 Evaluation Benchmarks and Metrics
We conduct experiments primarily on three benchmarks: Spot-the-Diff [12], Image-Editing-Request [27], and Birds-to-Words [9], which are described in further detail below. Consistent with prior research, we evaluate our models using BLEU (B) [19], METEOR (M) [4], CIDEr (C) [34], and ROUGE-L (R) [16] as metrics. Specifically, Spot-the-Diff [12] dataset is the inaugural dataset designed for the task of image difference captioning. It comprises 13,192 precisely aligned images sourced from surveillance cameras, characterized by a consistent viewpoint. Each image pair is annotated with a concise description of all significant differences, averaging 1.86 captions per pair. Image-Editing-Request [27] dataset, sourced from 3,939 real image pairs derived from online image-editing forums, includes 5,695 editing directives. Typically, the modified objects in these images are subtle and indistinct. Birds-to-Words [9] dataset documents the intricate distinctions among various bird species gathered in natural settings. Annotators focus solely on the visual differences between birds without referencing specific bird categories or the environmental background. This approach ensures that annotations are both precise and relevant to the observed phenotypic variations.
Table 1: Comparison to SOTA on three IDC benchmarks based on BLEU-4 (B), CIDEr (C), METEOR (M), and ROUGE-L (R).
(a) Comparison on Image-Editing-Request.
@@ B C M R
OneDiff (ours) 29.6 109.6 25.1 55.6
(b) Comparison on Spot-the-Diff.
@@
(c) Comparison on Birds-to-Words.
@@
OneDiff (ours) 27.1 28.4 25.8 49.6
@@ 5.3 Performance Comparison
As shown in Tab. 1(a), we evaluate the performance of our model on Image-Editing-Request. OneDiff demonstrates significant superiority in this benchmark. Specifically, compared to the SOTA models, the improvement of our OneDiff on CIDEr is nearly doubled, i.e. 211% compare with SCORER+CBR. We then compare the performance on Spot-the-Diff. As detailed in Tab. 1(b), OneDiff achieves significant improvements across BLEU, CIDEr and METEOR compared with SOTAs. Finally, we analyze the performance disparities on the Birds-to-Words. Consistent with the previous two benchmarks, our model demonstrated a significant lead as shown in Tab. 1(c). Quantitative results comparing to other IDC methods are displayed in in Fig. 5. OneDiff generates more reasonable difference captions with correct details on the bad cases of these methods.
Figure 5: Examples of OneDiff generated results.
Table 2: Ablation of model design, including Coupled Sample Training (CST) strategy and Visual Delta Module (VDM). We conduct ablation analysis on three aforementioned benchmarks.
@@ B C M R B C M R B C M R
× × 23.7 85.2 22.0 47.6 9.0 42.0 13.2 31.8 22.4 18.5 24.8 46.1
× 24.5 102.3 23.5 52.1 10.7 48.6 14.1 33.6 25.5 23.0 25.7 46.4
× 26.1 91.8 23.0 49.8 10.8 46.1 13.6 32.4 24.1 22.1 25.3 45.9
29.6 109.6 25.1 55.6 12.8 56.6 14.6 35.8 27.1 28.4 25.8 49.6
Table 3: Ablation of language models and Visual Delta Tokens on Birds-to-Words.
(a) A more efficient language model.
LLM B C M R
Vicuna-7B 25.0 21.8 25.9 46.7
Qwen2-1.5B 23.8 20.1 25.1 46.7
(b) Impact of dropping tokens on performance of CIDEr.
Used Delta Tokens 100% 50% 20% 0%
OneDiff (Vicuna-7B) 21.8 20.8 19.6 18.1
@@
Coll 22.3 90.1 22.0 51.5 8.9 40.7 13.8 31.8 23.7 20.8 24.4 46.9
Coll+Syn 26.5 94.4 24.0 51.5 9.0 42.0 14.3 32.8 25.5 21.8 25.5 46.7
Coll+Real 29.5 96.8 24.3 52.4 10.5 44.7 14.6 33.4 26.5 22.0 26.4 47.5
Coll+Syn+Real 29.6 109.6 25.1 55.6 12.8 56.6 14.6 35.8 27.1 28.4 25.8 49.6